Circular No. 5861

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA IN M31 R. R. Treffers, A. V. Filippenko, B. Leibundgut, Y. Paik, L. F. M.

Lee, and T. Matheson, University of California at Berkeley; and M. W. Richmond, Princeton University, report their discovery of a nova in M31 (α = 0^{h}40^{m}00^{s}, δ = +40^{o}59′.7, equinox 1950.0). The object was found during the Leuschner Observatory Supernova Search, which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera. It was detected in an image obtained on Aug. 27 UT, at R = 15.9 ± 0.3, located ∼ 10′′ east and 101′′ south of the nucleus (or 8′′ east and 14′′ south of a nearby star). Confirmation was made on Aug. 29 with the 0.50-m Berkeley Automatic Imaging Telescope, but the nova had already faded to R = 16.8 ± 0.3. No object appears at this position to limits of R = 16.3, 16.2, and 15.9 in images obtained on Aug. 18, 21, and 24, respectively. A noisy spectrum (range 310-1000 nm) obtained on Sept. 12 with the Shane 3-m reflector at Lick Observatory reveals unusually broad Hα and Hβ emission lines (FWHM ∼ 4400 km/s) superposed on a weak continuum.

GALACTIC CENTER R. Sunyaev, E. Churazov, B. Cordier, and M. Denis, on behalf of the

GRANAT team (Space Research Institute, Moscow; Service d’Astrophysique, Centre d’Etudes de Saclay; Centre d’Etude Spatiale des Rayonnements, Toulouse), report: “On Aug. 28, GRANAT started a new set of Galactic Center observations in hard x-rays. The observations were performed on, or are planned for, Aug. 28, 29, and 31, Sept. 1, 2, 4, 5, 6, 8, 9, 10, 12, 13, 14, 17, 18, 25, 26, 29, and 30, and Oct. 3 and 4 (and probably a few days more). Preliminary analysis revealed the following sources in 35to 100-keV images, averaged over the period Aug. 28-Sept. 3: 1E 1740.7– 2942 at a level of 100 ± 5.4 mCrab; GRS 1758–258, 53 ± 7.1 mCrab; GX 1+4, 48.7 ± 7.1 mCrab; GX 354–0, 34.5 ± 6.62 mCrab.”

YY HERCULIS Visual magnitude estimates (cf. IAUC 5805): June 16.85 UT, 11.6

(A. Boattini, Florence, Italy); 30.07, 11.7 (P. Schmeer, Bischmisheim, Germany); 30.96, 11.4 (E. Schweitzer, Strasbourg, France); Aug. 9.91, 11.7 (A. Pereira, Cabo da Roca, Portugal); Sept. 9.86, 11.6 (Pereira).

1993 September 13 Daniel W. E. Green